OpenStudy (anonymous):

what is the equation of a parabola with vertex (0,0) and focus (0,4)

OpenStudy (mathmate):

The focus is above the vertex, so the parabola opens up. The distance c=4 between focus and the vertex is the same distance as between vertex and directrix. The equation of the parabola is \(x^2=4cy\) where c=4 so the equation of the parabola is \(\huge y=\frac{x^2}{4c} \)

OpenStudy (mathmate):



OpenStudy (anonymous):

use this...put values and see.. http://www.endmemo.com/geometry/parabolaequation.php

OpenStudy (anonymous):

x^2=16y